x        3               
—————— = - ———
 2x-1       5

Answer by Edwin McCravy(20056)


   x        3               
—————— = - ———
 2x-1       5

The LCD = 5(2x-1), so put that over 1
and put

  5(2x-1)
 ————————
     1

to the right of both fractions:

   x    5(2x-1)       3   5(2x-1)               
——————·————————— = - ———·—————————
 2x-1      1          5      1
            
Cancel the (2x-1)'s on the left.
Cancel the 5's on the right

           1              1
   x    5(2x-1)       3   5(2x-1)               
——————·————————— = - ———·—————————
 2x-1      1          5      1
   1                  1

So we have 

              5x = -3(2x-1)

              5x = -6x + 3

             11x = 3

               x = 3/11
Edwin